Lura:

∠A and \angle B∠B are complementary angles. If m\angle A=(x-15)^{\circ}∠A=(x−15) ∘ and m\angle B=(3x-19)^{\circ}∠B=(3x−19) ∘ , then find the measure of \angle A∠A. plz help me i rlly need help!!!!

Lura:

1 attachment
TETSXPREME:

Well first you must find out what is x

Lura:

i tried and i keep getting like 20 and 18, but they're wrong-

TETSXPREME:

howd you get 20 and 18

Lura:

by doing diff formulas and saying they equal 180 and the other equals 90 degrees-

TETSXPREME:

Ok i think i found out what you did wrong

TETSXPREME:

you dont Ill give you the formula and you find the answer ok?

TETSXPREME:

4x-34=90 is the formula you solve

Lura:

4x=124?

TETSXPREME:

Yea and then divide

Lura:

31

TETSXPREME:

Try that

Lura:

no-

TETSXPREME:

IHHH OKK I GET IT NOW MBBB erase everything i said

Lura:

okiii

TETSXPREME:

Do 3x-19=90 since your looking for A you gotta find B

Lura:

1 attachment
TETSXPREME:

round it up so it would be 36

TETSXPREME:

90-36= ?

Lura:

54

surjithayer:

\[\angle A+\angle B=90\] x-15+3x-19=90 4x-34=90 4x=90+34=124 x=124/4=31 \[\angle A=x-15=31-15=16\]

Lura:

thx ^^

Lura:

∠A and \angle B∠B are vertical angles. If m\angle A=(6x-26)^{\circ}∠A=(6x−26) ∘ and m\angle B=(5x+6)^{\circ}∠B=(5x+6) ∘ , then find the measure of \angle B∠B.

Lura:

no- it said it was wrong-

TETSXPREME:

send screenshot

TETSXPREME:

@lura

Lura:

1 attachment
TETSXPREME:

do line b equals 180

TETSXPREME:

and then 180 minus your answer

TETSXPREME:

and that should give u line a

Lura:

bruh istf delta mathhhhh is so dumb it said not correct

surjithayer:

6x-26=5x+6 6x-5x=6+26 x=32 \[\angle B=5x+6=5\times 32+6=166\]